OpenStudy (anonymous):

Jerome ordered some CDs and DVDs online. He bought 2 CDs and 5 DVDs. He spent a total of $138.45. Part 1: Write the equation for this word problem. Be sure to identify your variables. Part 2: If each CD cost $14.25, solve the equation to find the cost of each DVD. Show your work.

OpenStudy (anonymous):

I am having trouble writing the equation for this problem Please help!!

OpenStudy (anonymous):

I am having trouble writing the equation for this problem Please help!!

OpenStudy (anonymous):

I am having trouble writing the equation for this problem Please help!!

OpenStudy (anonymous):

Let x = # of CDs Let y = # of DVDs so 2x + 5y = 138.45

OpenStudy (anonymous):

Does that make sense for the first part?

OpenStudy (anonymous):

Anybody home?

OpenStudy (anonymous):

Oh sorry my computer is slow so it's hard to respond fast

OpenStudy (anonymous):

Yes that makes perfect sense

OpenStudy (anonymous):

So if each cd costs $14.25 then I multiply it by 2 and I get $28.50 now I get the equation 28.50 + y = 138.45 I don't know what to do next please help

OpenStudy (anonymous):

I figured it out ranga thank you though

OpenStudy (anonymous):

My answer that I go t was $9.90 per DVD

OpenStudy (ranga):

The original equation was: 2x + 5y = 138.45 x is the price of each CD y is the price of each DVD In part 2 they tell you each CD costs $14.25 So x = 14.25 Put that in the equation above: 2(14.25) + 5y = 138.45 28.5 + 5y = 138.45 subtract 28.5 from both sides 5y = 109.95 divide both sides by 5 y = 21.99 Each DVD costs $21.99

OpenStudy (anonymous):

Thank You!! :)

OpenStudy (ranga):

You are welcome!